What is the kinetic energy of the continent?
Answer in units of J.
A jogger (of mass 60 kg) has the same kinetic
energy as that of the continent.
What would his speed be?
Answer in units of m/s.
1
Expert's answer
2020-12-21T11:01:43-0500
"K=0.5mv^2\\4140=0.5(60)v^2\\v=11.7\frac{m}{s}"